OpenStudy (anonymous):

Domain is (-\inf,5)U(5,\inf) what is the range and how do you find it? For function y=4/(5-x)

OpenStudy (turingtest):

y is not defined when the denominator is zero, because you cannot divide by zero

OpenStudy (anonymous):

\[so (-\infty,0)U(0,\infty)\]

OpenStudy (anonymous):

derivative y =4/(5-x) y' =4/(5-x)^2